40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Dec 29 21:30:25 2014, rana, Update, General, recovery 
    Reply  Mon Jan 5 18:15:18 2015, Jenne, Update, CDS, iscex reboot 
Message ID: 10846     Entry time: Mon Dec 29 21:30:25 2014     Reply to this: 10853
Author: rana 
Type: Update 
Category: General 
Subject: recovery 
  1. Control room is at +66 F. Brrrr.
  2. Alignment of input beam into the IMC was wacky; locked on HOM.
  3. Re-aligned beam into the PMC first.
  4. Restarted mxstream for c1sus.
  5. Power cycled Martian router; all laptops were lost. Now better.
  6. Aligned launch beam from PSL to get onto the MCWFS better, MC is locking OK now. Moved MC SUS a little to get back to OSEM values from 6 days ago.
  7. Fixed LOCK_MC screen quad displays to be cooler.
  8. changed many of the ezcawrite calls in the mcup / mcdown to be 'caput -l' for more robustness. Still need ezcawrite for the binbary calls.
  9.  I didn't touch the mirrors on the MC REFL path, so we can still use that as a reference once the temperature returns to normal; the PSL room temp is down to 20C from 22 C a couple days ago.
  10. TRX values coming in to the LSC were frozen and the TRY_OUT16 was going to huge values even though camera flashes were reasonable. Tried restarting c1lsc model. No luck.
  11. Also tried shutdown -r now on c1lsc. No luck. Probably needs a RFM boot.
  12. Increased the FSS SLOW servo threshold to 9999 counts to avoid it running on some misaligned TEM01 mode locks. Increased the PID's I gain from 0.05 to 0.356 by tuning on some step responses as usual.
  13. By midnight the control room temperature is back around 71 F.
ELOG V3.1.3-